OpenStudy (anonymous):

Find the interval [a,b] for which the value of the integral b s(2+x-x^2)dx a is maximum

OpenStudy (anonymous):

help me :'(

OpenStudy (unklerhaukus):

is this your question ?\[\int\limits_a^b(2+x-x^2)\text dx\]

OpenStudy (anonymous):

yes!!! sry first time here....

OpenStudy (anonymous):

is the answer [-1, 2] or [-1,1/2] ??

OpenStudy (helder_edwin):

as @UnkleRhaukus pointed out, you have maximize \[ \large F(a,b)=\int_a^b(2+x-x^2)\,dx \]

OpenStudy (helder_edwin):

which one gives u the greatest area?

OpenStudy (helder_edwin):

*value, sorry.

OpenStudy (anonymous):

doesnt [-1, 1/2] wait no... it's [-1, 2]

OpenStudy (unklerhaukus):

what did you get when you integrated

OpenStudy (anonymous):

27/6 ???

OpenStudy (helder_edwin):

which interval?

OpenStudy (anonymous):

[-1,2]

OpenStudy (unklerhaukus):

\[\int\limits_a^b(2x^0+x^1-x^2)\text dx=\frac{2x^1}1+\frac{x^2}2-\frac{x^3}{3}\Big|_a^b\]\[=2(b-a)+\frac{(b-a)^2}2-\frac{(b-a)^3}{3}\] right?

OpenStudy (anonymous):

yes!

OpenStudy (unklerhaukus):

actually i think i made a mistake

OpenStudy (anonymous):

it's separated right

OpenStudy (anonymous):

[2b+b^2/2-b^3/3]-[2a+2a^2/2-a^3/3]

OpenStudy (unklerhaukus):

\[\int\limits_a^b(2x^0+x^1-x^2)\text dx=\frac{2x^1}1+\frac{x^2}2-\frac{x^3}{3}\Big|_a^b\]\[\qquad\qquad=\left({2b}+\frac{b^2}2-\frac{b^3}3\right)-\left({2a}+\frac{a^2}2-\frac{a^3}3\right)\]yeahs

OpenStudy (unklerhaukus):

so when is this maximum

OpenStudy (anonymous):

at a= -1 b= 2 ! right?

OpenStudy (sirm3d):

Maximize F(b) by derivative, minimize F(a) by derivative.

OpenStudy (unklerhaukus):

\[=b\left({2}+\frac{b}2-\frac{b^2}3\right)+a\left(-{2}-\frac{a}2+\frac{a^2}3\right)\]

OpenStudy (unklerhaukus):

\[=6b\left(12+3b-2b^2\right)+6a\left(2a^2-3a-12\right)\]

OpenStudy (sirm3d):

\[\large F \prime (b)=2+b-b^2=0\]\[\large \text{max at } b=2\]\[\large F \prime (a)=2+a-a^2=0\]\[\large \text{max at } a=-1\] interval: [-1, 2]

OpenStudy (sirm3d):

wow, my answer is among the choices given.

OpenStudy (anonymous):

omg lol

OpenStudy (sirm3d):

typo mistake. it should be min at a = -1

OpenStudy (anonymous):

yayyyyyy tytytyty! :')